how do you factor these problems 3x+6y divided by x+2y. Its a rational expression
Found 2 solutions by MathLover1, rothauserc:
Answer by MathLover1(20849)
Answer by rothauserc(4718)

we are given
(3x+6y) / (x+2y) = 3*(x+2y) / (x+2y)
cancel numerator and denominator
3*(x+2y) / (x+2y) = 3